x+(2x+0.15)+(x+0.05)=1.8

Answer by farohw(175)


x + (2x+0.15) + (x+0.05) = 1.8

x + 2x + 0.15 + x + 0.05 = 1.8

Now add: 4x + 0.20 = 1.8

4x = 1.8 - 0.20 --> 4x = 1.6

Divide:
4x/4 = 1.6/4

x = 0.40